OpenStudy (kkutie7):

I can't seem to do this integral: \[\int xarcsin(x^{2})dx\]

random231 (random231):

take x^2=sin(theta)

OpenStudy (kkutie7):

I don't see how that would help anything

OpenStudy (kkutie7):

I was thinking about integration by parts because I know the derivative for arcsin but i dont know

random231 (random231):

hmmm yeah by parts would help too! even if you substitute x^2 as sin0 it would ultimately be solved by parts. so yeah do it

OpenStudy (kkutie7):

so, \[u=arcsin(x^{2})\rightarrow du=\frac{1}{\sqrt{1-(x^{2})^{2}}}dx\] \[dv=xdx\rightarrow v=\frac{x^{2}}{2}\] \[\frac{x^{2}}{2}arcsin(x^{2})-\frac{1}{2}\int \frac{x^{2}}{\sqrt{1-x^{4}}}dx\]

OpenStudy (kkutie7):

If I try to integrate by parts again I come back to the same answer as before.

OpenStudy (bibby):

how are you integrating that by parts?

OpenStudy (bibby):

It looks like a trig integral to me

OpenStudy (kkutie7):

it is a trig integral. i dont understand how to do it. so i broke it into part. arcsin and x

OpenStudy (kkutie7):

what about u substitution \[\frac{1}{2}\int arcsin(u)dx \] still I don't know the integral of arcsin

OpenStudy (bibby):

sorry, didn't notice this tab, okay let's see \(u=x^2\implies \large \int \dfrac{u}{\sqrt{1-u^2}}du\\ s=1-u^2\\ds=-2udu\implies -\dfrac{1}{2}ds=udu \\ \large -\dfrac{1}{2}\int \dfrac{1}{\sqrt{s}}ds=-\dfrac{1}{2}2\sqrt s=-\sqrt s= -\sqrt{1-u^2}=-\sqrt{1-x^4} \\\large \frac{x^{2}}{2}\arcsin(x^{2})-\frac{1}{2}\int \frac{x^{2}}{\sqrt{1-x^{4}}}dx=\frac{x^{2}}{2}\arcsin(x^{2})-\dfrac{1}{2}*-\sqrt{1-x^4} \\ \huge \dfrac{1}{2}({x^{2}}\arcsin(x^{2})+\sqrt{1-x^4}) \)

OpenStudy (irishboy123):

\(+ \; C\) lol!!!!

OpenStudy (kkutie7):

@bibby question \[u=x^{2}\implies \large \int \frac{u}{\sqrt{1-u^2}}du\] why not \[u=x^{2}\implies \large \frac{1}{2}\int \frac{u}{\sqrt{1-u^2}}\frac{du}{\sqrt{u}}\]